5.Suppose that 14 children, who were learning to ride two-wheel bikes, were surveyed to determine how long they had to use training wheels. It was revealed that they used them an average of nine months with a sample standard deviation of four months. Assume that the underlying population distribution is normal.
a. Construct a 99% confidence interval for the population mean length of time using training wheels.
(i) State the confidence interval
sketch the graph
calculate error bound
1
Expert's answer
2020-09-16T17:47:03-0400

Let  the random variable "X" be the time of using training wheelsin months

The provided sample mean is "\bar{X}=9" and the sample standard deviation is "s=4." The size of the sample is "n=14" and the required confidence level is 99%.

The number of degrees of freedom are "df=14-1=13," and the significance level is "\alpha=0.01."

Based on the provided information, the critical t-value for "\alpha=0.01" and "df=13"

degrees of freedom is "t_c=3.0123." The 95% confidence for the population "\mu"  is computed using the following expression


"CI=(\bar{X}-\dfrac{t_c\times s}{\sqrt{n}}, \bar{X}+\dfrac{t_c\times s}{\sqrt{n}})="

"=(9-\dfrac{3.0123\times 4}{\sqrt{14}}, 9+\dfrac{3.0123\times 4}{\sqrt{14}})="

"=(5.78, 12.22)"

"EBM=\dfrac{t_c\times s}{\sqrt{n}}=\dfrac{3.0123\times 4}{\sqrt{14}}=3.22"